OpenStudy (anonymous):

Identify the number of solutions: -6x+6y=-4 2x-2y=5

OpenStudy (anonymous):

Please show work(: I need help on this

OpenStudy (anonymous):

if you divide the first equation by -3, notice the left side of both equations will look exactly alike but their right sides are different... do you know what that means?

OpenStudy (anonymous):

infinite many solutions because the determinent of the system is zero | -6 6 2 -2|=12-12=0

OpenStudy (anonymous):

@byteMe no,no that much honestly

OpenStudy (anonymous):

it means that they both have the same slope, BUT different y-intercepts... these two lines don't intersect so there is no solution to the system. these two lines are parallel...

OpenStudy (anonymous):

so then that will make it no solution right(also 0)??

OpenStudy (anonymous):

correct... no solution....

OpenStudy (anonymous):

@sami-21 no,what your saying is many solutions

OpenStudy (anonymous):

infinite solutions means they're the same line, or the lines are on top of each other.... in that case, each x, y that work for 1 line also works for the other... see if that works in this problem....

OpenStudy (anonymous):

wait,before i do that.aint im suppose to put it in y intercept form first.that's what my teacher said

OpenStudy (anonymous):

yes you can... that's a good way of "seeing" each lines slope and y-intercept....

OpenStudy (anonymous):

so for the first line i got y=1x-2/3 is that right? i did it in y intercept form

OpenStudy (anonymous):

yes..